OpenStudy (anonymous):

Which is the equation of a line that passes through the point (3, -4) and has a slope of 2. Select one: a. y = 2x - 10 b. y = 2x - 11 c. y = 2x - 4 d. y = 2x - 2

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 3}}\quad ,&{\color{blue}{ -4}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= 2 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (mrnood):

the standard straight line equation is y=mx+b In th at equation 'm' represents the slope. The question tells you that the slope is 2 so your equation is y=2x+b BUT you know that the point (3,-4) lies on the line so when y=-4 you know that x= 3 so -4=2*3 +b Solve for b = and you have oyur answer

OpenStudy (anonymous):

is the answer C?

OpenStudy (mrnood):

@Twix47 Is that a guess? Just solve the last equation for the value of b and you have your answer -4=2*3 +b

OpenStudy (anonymous):

@MrNood no i was asking if i was right?

OpenStudy (mrnood):

if you are not right - are you then going to go 'is it d' 'is it a' 'is it b' Try to solve the equation by following the help given above.

OpenStudy (anonymous):

no im not dang but 2*3=6 so it would then look like -4=6+b so actually wouldnt it be D because 6-4=2 so it would look like y=2x-2

OpenStudy (mrnood):

so far you ar correct that 2*3 =6 SO the equation is -4= 6+b You need to solve that correctly for the value of b

OpenStudy (anonymous):

but what is b?

OpenStudy (mrnood):

the equation for any straight line is y=mx+b m is the slope (which oyu know is 2) so your answer is in the form y=2x+b BUT you know that the point ( 3,-4) is on the line SO when y=-4 you know x=3 SO -4=2*3+b Work out what b is from the above equation then you have your answer: y=2x+b

OpenStudy (anonymous):

i got y=2x-4

OpenStudy (mrnood):

you know that when y=-4 then x = 3 Does your equation work for that point? i.e. does -4=2*3 - 4 If not then your answer is wrong

OpenStudy (mrnood):

PLEASE try to solve this equation -4=2*3+b Work out what b is from the above equation then you have your answer: y=2x+b

OpenStudy (anonymous):

-4=6-4 so it would be D y=2x-2

OpenStudy (mrnood):

you know that when y=-4 then x = 3 Does your equation work for that point? i.e. does -4=2*3 - 2 If not then your answer is wrong